dr.kawasaki is twice as old as his son. twenty years ago, he was 4 times as old as his son was ten. how old are they now?
Answer by mananth(16949)

son----x years
Kawasaki ----- 2x
20 years ago
son = x-20
Kawasaki = 2x-20
..
2x-20= 4(x-20)
2x-20=4x-80
2x=60
x= 30 son's age
father's age = 30*2=60 years